OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x - 7 y = - one halfx - 2 y = 2x - 2 y = - one halfx - 7

OpenStudy (nurali):

The slope of the given equation: y=2x+2 So, the slope of the new equation is the same: 2 Use the formula: y=m(x)+b and input the slope and given point: 3=2(5)+b Solve for b: 3=10+b b=-7 Use the same formula, without the (y) and (x) values, use the variables: y=2x-7 Answer: y=-2x+14 -OR- 2x+y=14